I also have another question, here it is. It is on the substitution method. Here it is 3b-2a=4 and b/2-2a/3=1
Answer by uma(370)


3b - 2a = 4
==> 3b = 4 + 2a [adding 2a to both the sides]
==> b = (4+2a)/3
Plugging in this value in the second equation,
b/2 - 2a/3 = 1
==> (4+2a)/3*2 - 2a/3 = 1
==> 2(2+a)/6 - 2a/3 = 1 [factoring 4+2a]
==> 2+a/3 - 2a/3 = 1 [cancelling 2 in the first term]
==> 2 + a - 2a = 3 [multiply by 3]
==> 2 - a = 3
==> -a = 3 - 2
==> -a = 1
==> a = -1
so b = 4+2(-1)/3
==> b = 2/3
Thus a = -1 amd b= 2/3


Good luck